If Sally bought three chocolate bars and a pack of gum and paid $1.75. Jake bought two chocolate bars and four packs of gum and paid $2.00. How will I find the cost of a chocolate bar and the cost of a pack of gum
Found 2 solutions by josmiceli, Edwin McCravy:
Answer by josmiceli(19441)

c = cost of chocolate bar
g = cost of a pack of gum
3c + g = 1.75
2c + 4g = 2.00
multiply 1st by 4
and subtract the 2nd
12c + 4g = 7.00
-2c - 4g = -2.00
10c = 5.00
c = .50
so, cost of a chocolate bar is $.50
2(.50) + 4g = 2.00
1.00 + 4g = 2.00
4g = 1.00
g = .25
so, cost of pack of gum is $.25
check
3(.50) + .25 = 1.75
1.50 + .25 = 1.75
OK

Answer by Edwin McCravy(20086)

If Sally bought three chocolate bars and a pack of gum and paid $1.75. Jake
bought two chocolate bars and four packs of gum and paid $2.00. How will I find
the cost of a chocolate bar and the cost of a pack of gum.

Let the cost of a chocolate bar be C¢.
Then the cost of three chocolate bars was 3C¢.
The cost of two chocolate bars was 2C¢.

Let the cost of a pack of gum be G¢.
Then the cost of four packs of gum was 4G¢. 

>>...Sally bought three chocolate bars and a pack of gum and paid $1.75...<<

Translation:

The cost of three chocolate bars plus the cost 
of a pack of gum was 175¢.

                   3C + G = 175

>>...Jake bought two chocolate bars and four 
packs of gum and paid $2.00...<<

Translation:

The cost of two chocolate bars plus the cost 
of four packs of gum was 200¢.

                 2C + 4G = 200

So we have this system of equations:

                 3C +  G = 175
                 2C + 4G = 200

Can you solve that system?  Solve the first 
for G, then substitute in the second to find 
C, then substitute that in the first to find 
G.  If you can't do that, then post again 
asking how to do that.

Answer:  A chocolate bar costs 50¢,
         A pack of gum costs 25¢.

Edwin